# Little Red Dots from Ultra-Strongly Self-Interacting Dark Matter

### v1

@@ Abstract
We investigate the possibility that the recently identified population of high-redshift, obscured quasars—known as “Little Red Dots” (LRDs)—originates from early black hole seed formation driven by ultra-strongly self-interacting dark matter (uSIDM). In this framework, dark matter halos undergo gravothermal core collapse due to large self-interaction cross sections, resulting in the rapid formation of massive black hole (BH) seeds with masses at redshifts . We develop a semi-analytic model that tracks the evolution of the dark matter halo population, the redshift of collapse , and the corresponding BH mass function. Black hole growth is modeled stochastically via a log-normal Eddington ratio distribution and a finite duty cycle. We find that the uSIDM scenario naturally reproduces key observed properties of LRDs, including their abundance, compactness, and characteristic BH masses, while offering a mechanism for early, obscured black hole formation that is difficult to achieve in standard CDM-based models. The predicted SMBH mass function at shows excellent agreement with LRD observational data and SIDM merger-tree simulations, particularly at the high-mass end (). These results suggest that LRDs may serve as powerful observational tracers of exotic dark sector physics and that SMBH formation in the early universe could be significantly shaped by non-gravitational dark matter interactions.
@@ 1 Introduction
The discovery of “Little Red Dots” (LRDs) through deep infrared observations with the James Webb Space Telescope (JWST) has revolutionized our understanding of supermassive black hole (SMBH) formation and growth in the early Universe. These enigmatic objects represent a population of compact, heavily dust-reddened active galactic nuclei (AGN) observed at high redshifts (), characterized by their distinctive red rest-frame optical emission coupled with blue ultraviolet (UV) slopes [1, 2, 3]. The emergence of LRDs as a distinct class of high-redshift sources has fundamentally challenged existing models of early quasar evolution and SMBH assembly, necessitating a comprehensive reevaluation of the mechanisms driving black hole growth during the first billion years of cosmic history.
@@
Theoretical studies reinforce these observational constraints. Kido et al. explore the consequences of embedding a super-Eddington disc within a massive optically thick envelope to avoid disruptive outflows [9]. Although this envelope could stabilize the system, it is highly contrived and fails to naturally reproduce the observed spectral energy distributions. Meanwhile, alternative scenarios invoking sub-Eddington fueling of primordial black hole seeds offer a more straightforward match to LRD abundance, luminosity, and red colors [10]. In particular, the sub-Eddington accretion onto supermassive primordial seeds naturally accounts for the X-ray faintness, compact sizes, and moderate luminosities without requiring finely tuned obscuration or envelopes. Thus, both empirically and theoretically, super-Eddington models are increasingly disfavored in favor of simpler, sub-Eddington interpretations.
@@
The complex dust geometry and reddening mechanisms observed in LRDs present a fascinating interplay between AGN physics and circumgalactic medium dynamics. The primary source of observed reddening is attributed to AGN-driven dusty outflows, which create extended distributions of obscuring material characterized by substantial dust covering fractions and complex three-dimensional structures around the central black hole [11, 12]. These outflow-dominated systems exhibit distinctive spectral properties that require sophisticated modeling approaches, often invoking clumpy or anisotropic dust distributions combined with non-standard attenuation laws to adequately reproduce their observed spectral energy distributions (SEDs). This complexity distinguishes LRDs from conventional dusty starburst galaxies and highlights the unique physical conditions present in these extreme accretion environments [1, 6].
@@
The remarkable properties of LRDs—particularly their rapid SMBH growth rates and unexpectedly high abundance—have motivated the exploration of alternative formation scenarios that extend beyond conventional astrophysical mechanisms. One particularly intriguing possibility involves the role of exotic dark matter physics, specifically self-interacting dark matter (SIDM) and its potential connection to the observed LRD population. Unlike standard cold dark matter, SIDM allows dark matter particles to undergo significant collisional interactions, fundamentally altering the internal structure and evolution of dark matter halos [16, 17]. In sufficiently dense environments, these interactions can drive a phenomenon known as core collapse, where energy redistribution within the halo leads to a runaway increase in central density that could profoundly influence the formation and growth of embedded SMBHs.
The core collapse mechanism in SIDM halos presents a compelling framework for understanding the rapid assembly of massive black hole seeds in the early Universe [18, 19]. During the collapse process, the dramatic concentration of dark matter in halo centers creates gravitational potential wells capable of efficiently funneling baryonic material into compact regions, potentially facilitating the formation of intermediate-mass black hole seeds with masses on the order of – [20]. These substantial seed masses could provide a significant head start for subsequent accretion processes, offering a natural explanation for how SMBHs with masses approaching could assemble within the limited timescales available in the first billion years after the Big Bang. This scenario addresses one of the most pressing challenges in early Universe SMBH formation: the difficulty of growing black holes from stellar-mass seeds to supermassive scales through standard accretion processes alone within the available cosmic time.
The SIDM framework also provides a potential explanation for several puzzling observational characteristics of LRDs. The high abundance of these systems at –, where they outnumber unobscured quasars by orders of magnitude, could reflect the prevalence of SIDM core collapse events during epochs when dark matter densities and halo interaction rates were conducive to such phenomena. Furthermore, the remarkably compact nature of LRDs (typically pc) might be a natural consequence of baryon concentration within collapsed dark matter cores, where the enhanced gravitational focusing creates the dense, heavily obscured environments characteristic of these systems [21, 13]. The extreme accretion rates inferred for LRD black holes could similarly emerge from the efficient matter transport enabled by the steep gravitational potential gradients present in core collapsed halos.
@@ 2.2 Core Collapse Dynamics and Black Hole Seed Formation
Alternatively, and perhaps more plausibly from an observational standpoint, the uSIDM core collapse can serve as a catalyst for the formation of massive Population III star clusters or direct collapse black holes by creating the extreme gravitational environments necessary for such processes [41, 42]. In this scenario, the collapsed dark matter core provides the initial gravitational seed that enables the rapid accumulation of baryonic material into a compact, high-density region where conventional astrophysical processes can take over. The enhanced gravitational potential can overcome the typical obstacles to massive black hole formation, such as stellar feedback from Population III supernovae or radiation pressure from accreting protoblack holes, by providing sufficiently deep potential wells to retain gas despite these disruptive processes.
@@ 2.3 Environmental Conditions and Halo Mass Dependence
The redshift dependence of uSIDM core collapse also plays a crucial role in determining the epoch of massive black hole formation. At higher redshifts, the enhanced cosmic density and reduced halo velocity dispersions create conditions that are particularly favorable for driving gravothermal instabilities [17, 47]. Thus, the number of uSIDM produced high-redshift quasars is roughly fixed after the collapse redshift. This redshift dependence naturally explains the observed prevalence of high-redshift quasars and as the environmental conditions become less conducive to uSIDM core collapse in more recent cosmic epoch, this provides a physical mechanism for the apparent decline in quasar space density at lower redshifts (roughly fixed number, but increasing spacial volume).
@@ 2.4 Observational Signatures and Constraints
The formation timescales and efficiency of uSIDM core collapse also predict specific correlations between quasar properties and their large-scale environments. Quasars formed through this mechanism should show enhanced clustering on intermediate scales, reflecting the preferential formation of massive black holes in regions of enhanced dark matter density where core collapse is most efficient [52, 53]. Additionally, the mass function of high-redshift quasars should exhibit characteristic features related to the halo mass dependence of the core collapse process, potentially providing a distinctive signature that could distinguish uSIDM formation from conventional astrophysical mechanisms.
@@ 2.5 Implications for Little Red Dots and Early Quasar Evolution
The uSIDM framework provides a compelling theoretical foundation for understanding the formation and properties of Little Red Dots as a distinct population of high-redshift quasars. The rapid assembly of massive black hole seeds through core collapse naturally explains the ability of LRDs to reach supermassive scales within the limited time available at , addressing one of the primary challenges posed by these systems [1, 7]. Moreover, the enhanced accretion rates enabled by the deep potential wells created during core collapse could account for the extreme luminosities and rapid variability observed in many LRD systems.
@@ 3.1 The uSIDM Model Equations
The predictions we present for the SMBH mass function rely on the uSIDM seeding mechanism first developed in detail in our previous work [59]. In this framework, a subdominant component of the dark matter—characterized by a large self-interaction cross section—undergoes gravothermal collapse within dark matter halos at high redshift. The onset of collapse is governed by the condition
@@ 3.2 Results
In Fig. 1, we compare our uSIDM-based predictions (colored points) with two reference datasets: the LRD mass function measurements from Kokorev et al. [61] (red diamonds), and SIDM-based simulations from Jiang et al. [18] (red shaded region). The three uSIDM curves correspond to different combinations of cross-section and cooling parameters ( and ), taken from the parameter space explored in Roberts et al. [59]. Specifically we use, , which are the green, blue, and purple dots respectively. We also plot the Monte Carlo Poisson error in each bin. Each uSIDM model successfully tracks the steepening of the BHMF at lower masses (), where the SIDM model begins to underproduce BHs relative to LRD observations. We also note that the SIDM simulations from Jiang et al. [18] assume that all of their BHs are active, so one should take the red shaded regions as an upper limit. We also mention that in ref. [18], they do not plot the error bars for the LRD BH mass estimates with the data points they show - they assume that the Eddington ratio is 1. The LRD points therefore represent a lower limit on the black hole masses, and in our Fig. 1, we restore the error bars from ref. [61].
@@ 4 Discussion
However, connecting the observed properties of LRDs to underlying SIDM physics requires careful consideration of the complex interplay between dark matter dynamics and baryonic processes. While the weak or undetected X-ray emission characteristic of many LRDs [15, 11] could potentially reflect the extreme obscuration expected in core collapsed environments, current observations primarily attribute such properties to AGN-driven dusty outflows and associated feedback mechanisms [1, 12]. The challenge lies in identifying observational diagnostics that can definitively distinguish between SIDM-influenced formation scenarios and purely baryonic processes, as both could potentially produce systems with similar electromagnetic signatures.
@@ 5 Conclusion
Beyond mass functions, the uSIDM hypothesis makes several testable predictions. LRD host galaxies should preferentially reside in high-density environments conducive to early core collapse, exhibit compact stellar and gas distributions (on scales ), and show signs of enhanced baryon inflow consistent with deep potential wells. Future observational programs—particularly those combining JWST spectroscopy, ALMA kinematic mapping, and large-scale structure clustering—could detect these signatures and discriminate between uSIDM and standard seeding pathways.
@@ Acknowledgments
This work is partly supported by the U.S. Department of Energy grant number de-sc0010107 (SP).

### v2

@@ Abstract
We investigate the possibility that the recently identified population of high-redshift, obscured quasars - known as “Little Red Dots” (LRDs) - originates from early black hole seed formation driven by ultra-strongly self-interacting dark matter (uSIDM). In this framework, dark matter halos undergo gravothermal core collapse due to large self-interaction cross sections, resulting in the rapid formation of massive black hole (BH) seeds with masses at redshifts . We develop a semi-analytic model that tracks the evolution of the dark matter halo population, the redshift of collapse , and the corresponding BH mass function. Black hole growth is modeled stochastically via a log-normal Eddington ratio distribution and a finite duty cycle. We find that the uSIDM scenario naturally reproduces key observed properties of LRDs, including their abundance, compactness, and characteristic BH masses, while offering a mechanism for early, obscured black hole formation that is difficult to achieve in standard CDM-based models. The predicted SMBH mass function at shows excellent agreement with LRD observational data and SIDM merger-tree simulations, particularly at the high-mass end (). These results suggest that LRDs may serve as powerful observational tracers of exotic dark sector physics and that SMBH formation in the early universe could be significantly shaped by non-gravitational dark matter interactions.
@@ 1 Introduction
The discovery of “Little Red Dots” (LRDs) through deep infrared observations with the James Webb Space Telescope (JWST) has revolutionized our understanding of supermassive black hole (SMBH) formation and growth in the early Universe. These enigmatic objects represent a population of compact, heavily dust-reddened active galactic nuclei (AGN) observed at high redshifts (), characterized by their distinctive red rest-frame optical emission coupled with blue ultraviolet (UV) slopes [1, 2]. The emergence of LRDs as a distinct class of high-redshift sources has fundamentally challenged existing models of early quasar evolution and SMBH assembly, necessitating a comprehensive reevaluation of the mechanisms driving black hole growth during the first billion years of cosmic history.
@@
Theoretical studies reinforce these observational constraints. Kido et al. [8] explore the consequences of embedding a super-Eddington disc within a massive optically thick envelope to avoid disruptive outflows [8]. Although this envelope could stabilize the system, it is highly contrived and fails to naturally reproduce the observed spectral energy distributions. Meanwhile, alternative scenarios invoking sub-Eddington fueling of primordial black hole seeds offer a more straightforward match to LRD abundance, luminosity, and red colors [9]. In particular, the sub-Eddington accretion onto supermassive primordial seeds naturally accounts for the X-ray faintness, compact sizes, and moderate luminosities without requiring finely tuned obscuration or envelopes. Thus, both empirically and theoretically, super-Eddington models are increasingly disfavored in favor of simpler, sub-Eddington interpretations.
@@
The complex dust geometry and reddening mechanisms observed in LRDs present a fascinating interplay between AGN physics and circumgalactic medium dynamics. The primary source of observed reddening is attributed to AGN-driven dusty outflows, which create extended distributions of obscuring material characterized by substantial dust covering fractions and complex three-dimensional structures around the central black hole [10, 11]. These outflow-dominated systems exhibit distinctive spectral properties that require sophisticated modeling approaches, often invoking clumpy or anisotropic dust distributions combined with non-standard attenuation laws to adequately reproduce their observed spectral energy distributions (SEDs). This complexity distinguishes LRDs from conventional dusty starburst galaxies and highlights the unique physical conditions present in these extreme accretion environments [5].
@@
The remarkable properties of LRDs - particularly their rapid SMBH growth rates and unexpectedly high abundance - have motivated the exploration of alternative formation scenarios that extend beyond conventional astrophysical mechanisms. One particularly intriguing possibility involves the role of exotic dark matter physics, specifically self-interacting dark matter (SIDM) and its potential connection to the observed LRD population. Unlike standard cold dark matter, SIDM allows dark matter particles to undergo significant collisional interactions, fundamentally altering the internal structure and evolution of dark matter halos [15, 16]. In sufficiently dense environments, these interactions can drive a phenomenon known as core collapse, where energy redistribution within the halo leads to a runaway increase in central density that could profoundly influence the formation and growth of embedded SMBHs.
The core collapse mechanism in SIDM halos presents a compelling framework for understanding the rapid assembly of massive black hole seeds in the early Universe [17, 18, 19]. During the collapse process, the dramatic concentration of dark matter in halo centers creates gravitational potential wells capable of efficiently funneling baryonic material into compact regions, potentially facilitating the formation of intermediate-mass black hole seeds with masses on the order of – [9]. These substantial seed masses could provide a significant head start for subsequent accretion processes, offering a natural explanation for how SMBHs with masses approaching could assemble within the limited timescales available in the first billion years after the Big Bang. This scenario addresses one of the most pressing challenges in early Universe SMBH formation: the difficulty of growing black holes from stellar-mass seeds to supermassive scales through standard accretion processes alone within the available cosmic time.
The SIDM framework also provides a potential explanation for several puzzling observational characteristics of LRDs. The high abundance of these systems at , where they outnumber unobscured quasars by orders of magnitude, could reflect the prevalence of SIDM core collapse events during epochs when dark matter densities and halo interaction rates were conducive to such phenomena. Furthermore, the remarkably compact nature of LRDs (typically pc) might be a natural consequence of baryon concentration within collapsed dark matter cores, where the enhanced gravitational focusing creates the dense, heavily obscured environments characteristic of these systems [12]. The extreme accretion rates inferred for LRD black holes could similarly emerge from the efficient matter transport enabled by the steep gravitational potential gradients present in core collapsed halos.
@@ 2.2 Core Collapse Dynamics and Black Hole Seed Formation
Alternatively, and perhaps more plausibly from an observational standpoint, the uSIDM core collapse can serve as a catalyst for the formation of massive Population III star clusters or direct collapse black holes by creating the extreme gravitational environments necessary for such processes [39, 40]. In this scenario, the collapsed dark matter core provides the initial gravitational seed that enables the rapid accumulation of baryonic material into a compact, high-density region where conventional astrophysical processes can take over[41]. The enhanced gravitational potential can overcome the typical obstacles to massive black hole formation, such as stellar feedback from Population III supernovae or radiation pressure from accreting protoblack holes, by providing sufficiently deep potential wells to retain gas despite these disruptive processes.
@@ 2.3 Environmental Conditions and Halo Mass Dependence
The redshift dependence of uSIDM core collapse also plays a crucial role in determining the epoch of massive black hole formation. At higher redshifts, the enhanced cosmic density and reduced halo velocity dispersions create conditions that are particularly favorable for driving gravothermal instabilities [16, 33, 34]. Thus, the number of uSIDM produced high-redshift quasars is roughly fixed after the collapse redshift. This redshift dependence naturally explains the observed prevalence of high-redshift quasars and as the environmental conditions become less conducive to uSIDM core collapse in more recent cosmic epoch, this provides a physical mechanism for the apparent decline in quasar space density at lower redshifts (roughly fixed number, but increasing spacial volume).
@@ 2.4 Observational Signatures and Constraints
The formation timescales and efficiency of uSIDM core collapse also predict specific correlations between quasar properties and their large-scale environments. Quasars formed through this mechanism should show enhanced clustering on intermediate scales, reflecting the preferential formation of massive black holes in regions of enhanced dark matter density where core collapse is most efficient [49]. Additionally, the mass function of high-redshift quasars should exhibit characteristic features related to the halo mass dependence of the core collapse process, potentially providing a distinctive signature that could distinguish uSIDM formation from conventional astrophysical mechanisms.
@@ 2.5 Implications for Little Red Dots and Early Quasar Evolution
The uSIDM framework provides a compelling theoretical foundation for understanding the formation and properties of Little Red Dots as a distinct population of high-redshift quasars. The rapid assembly of massive black hole seeds through core collapse naturally explains the ability of LRDs to reach supermassive scales within the limited time available at , addressing one of the primary challenges posed by these systems [6]. Moreover, the enhanced accretion rates enabled by the deep potential wells created during core collapse could account for the extreme luminosities and rapid variability observed in many LRD systems.
@@ 3.1 The uSIDM Model Equations
The predictions we present for the SMBH mass function rely on the uSIDM seeding mechanism first developed in detail in our previous work [42]. In this framework, a subdominant component of the dark matter - characterized by a large self-interaction cross section - undergoes gravothermal collapse within dark matter halos at high redshift. The onset of collapse is governed by the condition
@@ 3.2 Results
In Fig. 1, we compare our uSIDM-based predictions (colored points) with two reference datasets: the LRD mass function measurements from Kokorev et al. [1] (red diamonds), and SIDM-based simulations from Jiang et al. [17] (red shaded region). The three uSIDM curves correspond to different combinations of cross-section and cooling parameters ( and ), taken from the parameter space explored in Roberts et al. [42]. Specifically we use, , which are the green, blue, and purple dots respectively. We also plot the Monte Carlo Poisson error in each bin. Each uSIDM model successfully tracks the steepening of the BHMF at lower masses (), where the SIDM model begins to underproduce BHs relative to LRD observations. We also note that the SIDM simulations from Jiang et al. [17] assume that all of their BHs are active, so one should take the red shaded regions as an upper limit. We also mention that in ref. [17], they do not plot the error bars for the LRD BH mass estimates with the data points they show - they assume that the Eddington ratio is 1. The LRD points therefore represent a lower limit on the black hole masses, and in our Fig. 1, we restore the error bars from ref. [1]. We also mention that in the companion paper to [17], Shen et al. [56], they use a dissipative SIDM model to reproduce the SMBH and LRD mass functions. Similar to [17], they also require a duty cycle of 1 to match the LRD population (see their Figure 2).
@@ 4 Discussion
However, connecting the observed properties of LRDs to underlying SIDM physics requires careful consideration of the complex interplay between dark matter dynamics and baryonic processes. While the weak or undetected X-ray emission characteristic of many LRDs [14, 10] could potentially reflect the extreme obscuration expected in core collapsed environments, current observations primarily attribute such properties to AGN-driven dusty outflows and associated feedback mechanisms [11]. The challenge lies in identifying observational diagnostics that can definitively distinguish between SIDM-influenced formation scenarios and purely baryonic processes, as both could potentially produce systems with similar electromagnetic signatures.
@@ 5 Conclusion
Beyond mass functions, the uSIDM hypothesis makes several testable predictions. LRD host galaxies should preferentially reside in high-density environments conducive to early core collapse, exhibit compact stellar and gas distributions (on scales ), and show signs of enhanced baryon inflow consistent with deep potential wells. Future observational programs - particularly those combining JWST spectroscopy, ALMA kinematic mapping, and large-scale structure clustering - could detect these signatures and discriminate between uSIDM and standard seeding pathways.
@@ Acknowledgments
This work is partly supported by the U.S. Department of Energy grant number de-sc0010107 (SP). This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST.